OpenStudy (anonymous):

Find the area bounded by the two curves: x = 100000(12sqrt{y}-1) x = 100000(12sqrt{y} -1)/(9sqrt{y}) The definite integral has value:?

OpenStudy (experimentx):

\( x = 100000(12 \sqrt{y}-1) \) \( x = 100000(12 \sqrt{y} -1)/(9 \sqrt{y}) \)

OpenStudy (experimentx):

try graphing first.